OpenStudy (anonymous):

Find the critical numbers of the function. g(x) = x^1/6 - x^5/6

myininaya (myininaya):

\[g'(x)=\frac{1}{6}*x^{\frac{1}{6}-1}-\frac{5}{6}*x^{\frac{5}{6}-1}=\frac{1}{6}(x^\frac{-5}{6}-5x^\frac{-1}{6})\] \[=\frac{1}{6}*(\frac{1}{x^\frac{5}{6}}-\frac{5}{x^\frac{1}{6}})=\frac{1}{6}*\frac{x^\frac{1}{6}-5x^\frac{5}{6}}{x^\frac{1}{6}*x^\frac{5}{6}}\] \[=\frac{1}{6}*x^\frac{1}{6}*\frac{1-5x^\frac{4}{6}}{x}\] \[=\frac{1}{6}*x^\frac{1}{6}*\frac{1-5x^\frac{2}{3}}{x}\]

myininaya (myininaya):

so we want to find when f'=0 and when f' dne

myininaya (myininaya):

f'=0 when 1-5x^\frac{2}{3}=0\] f' dne when x=0

myininaya (myininaya):

\[1-5x^\frac{2}{3}=0\]

myininaya (myininaya):

\[1=5x^\frac{2}{3}\] \[\frac{1}{5}=x^\frac{2}{3}\]

myininaya (myininaya):

\[x=(\frac{1}{5})^\frac{3}{2}\]

myininaya (myininaya):

\[x=\frac{1}{5^\frac{3}{2}}=\frac{1}{5\sqrt{5}}=\frac{1}{5\sqrt{5}}*\frac{\sqrt{5}}{\sqrt{5}}=\frac{\sqrt{5}}{25}\]

OpenStudy (anonymous):

it is asking for 2 critical points

myininaya (myininaya):

if you have x^2=9 this give you two solutions right?

OpenStudy (anonymous):

-3 and 3

myininaya (myininaya):

right so you you have \[x^\frac{2}{3}=\frac{1}{5}\] what does this give you?

myininaya (myininaya):

\[x=-\frac{\sqrt{5}}{25}, \frac{\sqrt{5}}{25}\]

myininaya (myininaya):

and you actually have 3 critical numbers including x=0

OpenStudy (anonymous):

-sqrt(5)/25 doesnt work!

myininaya (myininaya):

oh wait you are right because we have even roots in our problem

myininaya (myininaya):

so we only have the two and there no more critical numbers

myininaya (myininaya):

are there anymore questions?

OpenStudy (anonymous):

oh ok! it works with x=0. Can you help me another?

OpenStudy (anonymous):

Find the critical numbers of the function on the interval 0 ≤ θ < 2π. g(θ) = 4 θ - tan(θ)

myininaya (myininaya):

well remember we found x=0 way up there lol a critical number is finding where f'=0 and where f' dne

myininaya (myininaya):

so what do you think g' is?

myininaya (myininaya):

\[(4\theta)'=?\] \[(tanx)'=?\]

OpenStudy (anonymous):

tanx'=sex^2(x)

OpenStudy (anonymous):

the other one is 1

OpenStudy (anonymous):

sorry i mean 4

myininaya (myininaya):

right! so we have \[g'(\theta)=4-\sec^2(x)\]

OpenStudy (anonymous):

ok then!

OpenStudy (anonymous):

do i just input the numbers

myininaya (myininaya):

now to find critical numbers we need to set g'=0 and determine where g' dne( make sure these are numbers in the domain of g) so we have \[g'(\theta)=4-\frac{1}{\cos^2 \theta}=\frac{4\cos^2 \theta-1}{\cos^2 \theta}\] when is bottom 0?

myininaya (myininaya):

so cosx is 0 when x=pi/2 and 3pi/2 right?

myininaya (myininaya):

so those are two critical numbers we also have to find when f'=0

myininaya (myininaya):

4cos^2x-1=0 4cos^2x=1 cos^2x=1/4

myininaya (myininaya):

\[cosx=\sqrt{\frac{1}{4}}=\frac{1}{\sqrt{4}}=\frac{1}{2}\] so when does cosx=1/2 when x=?

myininaya (myininaya):

do you got it?

OpenStudy (anonymous):

60

myininaya (myininaya):

and where else?

myininaya (myininaya):

300 degrees or you can say 5pi/3 so we have four critical numbers

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2pi/3 and 4pi/3